The following data are from a completely randomized design. In the following calculations, use

α = 0.05.

Treatment
1
Treatment
2
Treatment
3
63 82 69
46 73 55
54 87 62
37 70 46

xj

50 78 58

sj2

123.33 62.00 96.67

a) Find the value of the test statistic. (Round your answer to two decimal places.)

Find the p-value. (Round your answer to three decimal places.)

p-value =

b)

Use Fisher's LSD procedure to determine which means are different.

Find the value of LSD. (Round your answer to two decimal places.)

LSD =
Answer #1
Applying one way ANOVA: (use excel: data: data analysis: one way ANOVA: select Array):
Source SS df MS F P value
treatment 1664.00 2 832.00 8.85 0.0075
error 846.00 9 94.00
Total 2510.00 11

a)

value of the test statistic F =8.85

p value =0.0075

b)

critical value of t with 0.05 level and N-k=9 degree of freedom= tN-k= 2.262
Fisher's (LSD) for group i and j =(tN-k)*(sp*√(1/ni+1/nj)   = 15.51
Difference Absolute Value Conclusion
x1-x2 28.00 significant difference
x1-x3 8.00 not significant difference
x2-x3 20.00 significant difference

treatment 1 and 2 as well treatment 2 and 3 are different